Karl said if I divide my dogs age by 4 and subtract 2. I'll get the same result as when I divide my dogs age by 8.
Write an equation and solve

Answer by mananth(16946)

let dog's age be x
divide by 4

= x/4
subtract 2
x/4 -2 = x/8
x/4-x/8= 2
8x-4x / 32 = 2
4x=64
x=16 the dogs age